Differentiate the following functions (i) If 𝑦 = 𝑒 βˆ’3𝑑 sin 4𝑑, prove that 𝑑 2𝑦 𝑑π‘₯ 2 + 6 𝑑𝑦 𝑑π‘₯ + 2𝑦 = 0. (ii) Given that sin(π‘₯ 2 + 𝑦) = 𝑦 2 (3π‘₯ + 1), show that 𝑑𝑦 𝑑π‘₯ = 2π‘₯ cos(π‘₯ 2 + 𝑦) βˆ’ 3𝑦 2 2𝑦(3π‘₯ + 1) βˆ’ cos(π‘₯ 2 + 𝑦


1
Expert's answer
2021-09-27T23:26:07-0400

(i) First, we know that "𝑦 = 𝑒^{−3𝑡} \sin{(4𝑡)}", then we proceed to derive


"\\ \frac{dy}{dt}=\frac{d}{dt}(𝑒^{−3𝑡}\sin{(4𝑡)})
\\ \frac{dy}{dt}=\sin{(4𝑡)}\frac{d(𝑒^{−3𝑡} )}{dx}+𝑒 ^{−3𝑡} \frac{d(\sin{(4𝑡)})}{dx}
\\ \frac{dy}{dt}=\sin{(4𝑡)}(-3𝑒^{−3𝑡})+𝑒^{−3𝑡}(4\cos{(4t)})
\\ \frac{dy}{dt}=-3𝑒^{−3𝑡}\sin{(4𝑡)}+4𝑒^{−3𝑡}\cos{(4t)}
\\ \frac{dy}{dt}=𝑒^{−3𝑡} \Big[ 4\cos{(4t)} -3\sin{(4𝑡)} \Big]"


After the first derivative we can work on the second one:


"\frac{d^2y}{dt^2}=\frac{d}{dt}\Big(𝑒^{−3𝑡} [ 4\cos{(4t)} -3\sin{(4𝑡)} ] \Big)
\\ \frac{d^2y}{dt^2}=𝑒^{−3𝑡} \frac{d}{dt}\Big( 4\cos{(4t)} -3\sin{(4𝑡)} \Big)+ ( 4\cos{(4t)} -3\sin{(4𝑡)} ) \frac{d}{dt}\Big( 𝑒^{−3𝑡}\Big)
\\ \frac{d^2y}{dt^2}=𝑒^{−3𝑡} \big( -16\sin{(4t)} -12\cos{(4𝑡)} \big)+ ( 4\cos{(4t)} -3\sin{(4𝑡)} )\big(-3𝑒^{−3𝑡}\big)
\\ \frac{d^2y}{dt^2}=𝑒^{−3𝑡} \big( -16\sin{(4t)} -12\cos{(4𝑡)} -12\cos{(4t)} +9\sin{(4𝑡)} \big)
\\ \frac{d^2y}{dt^2}=𝑒^{−3𝑡} \big( -7\sin{(4t)} -24\cos{(4𝑡)}\big)"


Now we have


"𝑦 = 𝑒^{−3𝑡} \sin{(4𝑡)}
\\ \frac{dy}{dt}=𝑒^{−3𝑡} \Big[ 4\cos{(4t)} -3\sin{(4𝑡)} \Big]
\\ \frac{d^2y}{dt^2}=-𝑒^{−3𝑡} \big( 7\sin{(4t)} +24\cos{(4𝑡)}\big)"


And we can substitute this on the DE, "\frac{d^2y}{dt^2}+6\frac{dy}{dt}+2y=0", to find if this is true for y:


"\\ 𝑒^{−3𝑡} \big( -7\sin{(4t)} -24\cos{(4𝑡)}\big)+6\Big(𝑒^{−3𝑡} \big[ 4\cos{(4t)} -3\sin{(4𝑡)} \big]\Big)+2 𝑒^{−3𝑡} \sin{(4𝑡)}=0
\\ 𝑒^{−3𝑡} \Bigg( -7\sin{(4t)} -24\cos{(4𝑡)}+ 24\cos{(4t)} -18\sin{(4𝑡)} +2\sin{(4𝑡)} \Bigg)=0"


"\\ \implies -𝑒^{−3𝑡} \Bigg( 23\sin{(4𝑡)} \Bigg) \neq 0"


That implies that "𝑦 = 𝑒^{−3𝑡} \sin{(4𝑡)}" is not a solution for the DE "\frac{d^2y}{dt^2}+6\frac{dy}{dt}+2y=0".


(ii) Now we have two expressions, A and B:


"\sin(x^2 + 𝑦) = y^2(3x+ 1)
\\ \implies A=B"


We proceed to differentiate both terms and we have to use implicit differentiation:


"A'=\cos(x^2 + y)(2x +\frac{dy}{dx})
\\ A'=2x\cos(x^2 + y) +\frac{dy}{dx}(\cos(x^2 + y))"


"B'= 3y^2+2y(3x+ 1)\frac{dy}{dx}
\\ B'= 3y^2+2y(3x+ 1)\frac{dy}{dx}"


Now we substitute to find "\frac{dy}{dx}" :


"A'=B'
\\2x\cos(x^2 + y) -3y^2= (2y(3x+ 1)-\cos(x^2 + y))\frac{dy}{dx}
\\ \text{ }
\\ \implies \cfrac{dy}{dx} = \cfrac{2x\cos(x^2 + y) -3y^2}{ 2y(3x+ 1)-\cos(x^2 + y)}"


That last result is consistent with the expected expression.




Reference:

  • Thomas, G. B., & Finney, R. L. (1961).Β Calculus. Addison-Wesley Publishing Company.